Solve the system of equations then classify the system.
y=-x-11
3x-2y=-18

Answer by ewatrrr(24785)

 
Hi
Solve the system of equations then classify the system.
y = -x-11
3x -2y = -18
3x -2(-x-11) = -18
3x + 2x+22 = -18
5x = -40
x = -8 and y = -3 (y = -(-8) - 11 = -3)
Ordered pair(-8,-3) the solution for this Independent system
CHECKING our Answer***
3x -2y = -18
3*-8 + 2*3 = -18
A consistent system has at least one solution.
An inconsistent system has no solutions.
An independent system has exactly one solution.
A dependent system has infinitely many solutions.